OpenStudy (anonymous):

Find the second derivate:

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x^2-1};\]

OpenStudy (anonymous):

Well, first you need to find the first derivative, and this will take some use of the chain rule.

OpenStudy (anonymous):

I found the first derivate:\[\frac{ 2x }{ 3\sqrt[3]{(x^2-1)^2}}\]

OpenStudy (anonymous):

i need someone to explain me the second derivate

OpenStudy (anonymous):

To find the second derivative, you just take the derivative of the first derivative according to all the same rules.

OpenStudy (anonymous):

i derivate "mot a mot"\[\frac{ 2(3\sqrt[3]{(x^2-1)^2})-2x(\frac{ 4(x^3-x) }{ \sqrt[3]{(x^2-1)^4} }) }{ 3\sqrt[3]{(x^2-1)^2} }\]

OpenStudy (anonymous):

but i need a simplified form

OpenStudy (anonymous):

well, one thing you can do to simplify it to note that: \[x^3 -x = x(x^2-1)\] and that: \[\sqrt[3]{(x^2-1)^4} = (x^2-1)^{4/3}\]

OpenStudy (anonymous):

make \(x^2-1=b\) and \(db=2x\) then f(x)=\(b^{1/3}\) f'(x)=\(1/3b^{-2/3}db\) f''(x)=\(-2/9b^{-5/3}db\) now just substitute

OpenStudy (anonymous):

Because db in f'(x) is technically a function of x, wouldn't you have to derive that as well when finding f''(x) including using the product rule if so?